# Intention Analysis Makes LLMs A Good Jailbreak Defender

### v2

@@ Abstract
Aligning large language models (LLMs) with human values, particularly in the face of stealthy and complex jailbreak attacks, presents a formidable challenge. In this study, we present a simple yet highly effective defense strategy, i.e., Intention Analysis ($\mathbb{IA}$). The principle behind this is to trigger LLMs’ inherent self-correct and improve ability through a two-stage process: 1) essential intention analysis, and 2) policy-aligned response. Notably, $\mathbb{IA}$ is an inference-only method, thus could enhance the safety of LLMs without compromising their helpfulness [^1] Improving the safety of LLM with training-required methods Ouyang et al. (2022); Touvron et al. (2023) always necessitates great effort to strike a delicate balance between safety and helpfulness.. Extensive experiments on SAP200 and DAN benchmarks across Vicuna, ChatGLM, MPT, DeepSeek, and GPT-3.5 show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -46.5% attack success rate) and maintain the general helpfulness. Encouragingly, with the help of our $\mathbb{IA}$, Vicuna-7b even outperforms GPT-3.5 in terms of attack success rate. Further analyses present some insights into how our method works. To facilitate reproducibility, we release our code and scripts at: https://github.com/alphadl/SafeLLM_with_IntentionAnalysis.
@@
[…]
@@ 1 Introduction
Figure 1: Performance of our method on different jailbreak datasets. Our $\mathbb{IA}$ 1) consistently reduces Attack Success Rate ($\downarrow$) against both prompt-level (SAP and DAN) and token-level (AdvBench) jailbreaks, 2) achieves remarkable safety improvements for both SFT (Vicuna-7B&13B) and RLHF (ChatGLM-6B) models.
Despite numerous strategies focusing on defending LLMs against these jailbreak attacks (Li et al., 2024; Jain et al., 2023; Chen et al., 2024), their efficacy markedly decreases when handling those with stealthy and complex intention (see experimental results in Table 3). This type of jailbreak attacks are typically hand-crafted prompts by using more instructions and employing specific techniques, such as creating a virtual scenario where LLMs are assigned the role of “DAN” (Do Anything Now, Liu et al., 2023). Such stealthy and complex intention poses a great challenge for LLMs to maintain safety when confronting these jailbreak attacks (Shen et al., 2023; Bao et al., 2023).
@@
    We introduce $\mathbb{IA}$, a new method that significantly enhances LLMs safety in the context of sophisticated jailbreak attacks through an intention analysis mechanism.
@@
    Empirically, our robust $\mathbb{IA}$ significantly and consistently reduces the harmfulness of LLM outputs, while maintaining the helpfulness, achieving new state-of-the-art performance on several benchmarks, e.g., AdvBench.
@@ 2 Related Work
Despite significant efforts to align LLMs with human preference (Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Korbak et al., 2023; Miao et al., 2024), adversarial attackers can still elicit harmful responses from LLMs by “jailbreak” attacks (Shen et al., 2023; Liu et al., 2023). Current jailbreak attack methods are primarily classified into two categories: prompt-level and token-level (Chao et al., 2023). Prompt-level jailbreak is typically hand-crafted through human ingenuity and is semantically understandable in general Shen et al. (2023). For token-level jailbreak, a representative work is to automatically fetch a transferable attack suffix through the Greedy Coordinate Gradient (GCG) algorithm which maximizes the probability of the language model generating an affirmative and unsafe response Zou et al. (2023).
In this work, we aim to enhance the safety of LLMs when facing prompt-level jailbreaks with stealthy and complex intentions.
Safety-Enhancing Defensive Mitigations
Recently, numerous methods have been developed to reduce LLMs’ harmful generations at the inference stage. A branch of them mainly concentrates on controlling the content that LLMs can see by pre-processing the user inputs, such as perplexity filtering (Alon and Kamfonas, 2023; Jain et al., 2023), paraphrasing (Jain et al., 2023) and re-tokenization (Cao et al., 2023; Jain et al., 2023). Another branch focuses on exploiting LLMs’ intrinsic capabilities of self-correction and improvement against jailbreak attacks, such as letting LLMs self-evaluate their outputs (Helbling et al., 2023) or reminding of safety in system mode with conventional decoding Xie et al. (2023) or contrastive decoding Zhong et al. (2024).
While these existing methods show effectiveness in preventing LLMs from generating unsafe responses, their efficacy significantly diminishes when confronted with sophisticated jailbreak attacks due to the ignorance of the underlying stealthy and malicious intention. In contrast, our method handles such jailbreak attacks by leveraging the intrinsic intent recognition capabilities of LLMs, which leads to a significant enhancement of LLMs’ safety.
@@ 3.2 $\mathbb{IA}$: Intention Analysis
To achieve the above goal, we introduce $\mathbb{IA}$, a zero-shot intention analysis mechanism, to guide LLMs to explicitly identify and understand the underlying intention of a user query before facilitate a final response. Specifically, we devise a two-stage intention analysis instruction to accomplish the whole process, as illustrated in Figure 2(b): (1) essential intention analysis and (2) policy-aligned response.
@@
Models Methods SAP200 Avg.($\downarrow$) $\bm{\Delta}\left(\downarrow\right)$
Fraud Polit. Porno. Race Relig. Suici. Terro. Viole.
ChatGLM-6B Vanilla 36.5 29.0 52.5 45.5 47.5 61.0 34.0 60.0 45.8 -39.68
Ours 5.00 3.00 8.50 5.50 11.5 10.5 0.50 4.50 6.12
@@
Ours 1.50 0.50 0.00 0.00 0.50 0.00 0.00 0.00 0.31
@@
Ours 1.00 0.00 2.00 0.00 1.00 5.50 0.00 0.00 1.12
@@
Ours 21.0 16.5 20.0 22.0 16.5 18.0 19.0 21.0 19.2
@@
Ours 3.00 0.00 5.50 0.50 1.00 2.00 0.0 0.50 1.56
@@
Ours 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Table 1: Comparison of our $\mathbb{IA}$ and vanilla in SAP200 (comprising eight distinct sensitive topics) in terms of ASR (%). The best average results are highlighted in bold.
Models Methods DAN Avg.($\downarrow$) $\bm{\Delta}\left(\downarrow\right)$
Anarc. Basic Oppos. Toxic Advan. Start. Guide. Virtu.
ChatGLM-6B Vanilla 25.4 26.2 32.8 48.7 28.3 19.4 14.7 36.6 29.0 -23.52
Ours 7.73 4.10 12.3 8.15 6.14 1.04 3.33 1.06 5.48
@@
Ours 0.00 9.74 1.02 3.59 5.15 1.02 2.05 0.52 2.89
[…]
Ours 1.03 0.52 0.51 1.02 0.51 0.00 1.03 0.51 0.64
[…]
Ours 5.13 6.15 6.15 5.13 6.15 3.08 4.62 6.67 5.38
[…]
Ours 0.51 0.00 1.54 1.54 23.6 0.00 0.00 3.08 3.78
[…]
Ours 0.00 0.00 2.05 0.00 1.02 1.03 1.02 0.00 0.64
Table 2: Comparison of our $\mathbb{IA}$ and vanilla in DAN (including eight jailbreak communities) in terms of ASR (%). The best average results are highlighted in bold.
@@ 4.1 Setup
We experiment on two main categories of jailbreak attacks, including two representative prompt-level jailbreak datasets (i.e. DAN (Shen et al., 2023) and SAP200 (Deng et al., 2023a)), and one popular token-level jailbreak dataset (i.e. AdvBench (Zou et al., 2023). Besides, to evaluate the effect of our $\mathbb{IA}$ on helpfulness for general in-distribution queries, we conduct experiments on three widely recognized datasets, i.e., AlpacaEval (Dubois et al., 2023), MMLU (Hendrycks et al., 2021) and TruthfulQA (Lin et al., 2022).
Methods AdvBench DAN SAP200
Vicuna-7B Vicuna-13B Vicuna-7B Vicuna-13B Vicuna-7B Vicuna-13B
Vanilla 83.0 87.0 48.4 60.0 70.0 65.9
+ BPE-dropout (Jain et al., 2023) 63.0 50.0 23.8 28.2 67.2 48.9
+ ICD (Wei et al., 2023) 1.00 0.00 44.4 58.9 32.8 32.8
+ Self Defense (Helbling et al., 2023) 24.0 20.0 31.3 28.8 53.2 29.7
+ Moral Self-Correction (Ganguli et al., 2023) 26.0 13.0 25.0 28.1 49.0 42.8
+ System-Mode Self-Reminder (Xie et al., 2023) 11.0 1.0 45.3 57.5 33.8 36.9
+ $\mathbb{IA}$ (Ours) 0.00 0.00 3.42 0.94 0.31 1.56
Table 3: Comparison of our method and existing advanced defense methods in terms of ASR (%). The best and second best results are highlighted in bold and underline.
Figure 3: The MultiJail Dataset results on ChatGPT under two jailbreak scenarios with and without our $\mathbb{IA}$. (a) results on direct MultiJail dataset. (b) results on malicious jailbreak prompt attached to MultiJail.
@@
For safety assessment, we annotate the harmfulness of responses and report attack success rate (ASR, Shen et al., 2023). Specifically, for SAP200 and AdvBench, we annotate harmfulness following Zou et al. (2023) by matching refusal strings (e.g., “I’m sorry”). For DAN, due to the complexity of responses, we adopt gpt-3.5-turbo-0613444https://openai.com/blog/chatgpt as the auto-annotation function following Deng et al. (2023a). See Appendix B for more detailed settings. In addition, we carry out human evaluations to ensure the credibility and accuracy of the assessment (See Section 5.3 for details).
For helpfulness assessment, we report win rate (Dubois et al., 2023) for AlpacaEval and accuracy (Hendrycks et al., 2021) for MMLU. For TruthfulQA, we follow Chuang et al. (2023) and report on two distinct metrics: MC1 and MC2 scores. For these helpfulness metrics, higher scores indicate stronger factuality.
@@
To evaluate the effectiveness of our $\mathbb{IA}$ method, we experiment on models varying scales and levels of alignment, including not only SFT models, i.e. Vicuna-7B/13B-v1.1 (Chiang et al., 2023) and MPT-30B-Chat (Team, 2023), but also RLHF models, i.e. ChatGLM-6B (Zeng et al., 2023) and DeepSeek-67B-Chat (DeepSeek-AI, 2024). Beyond these representative open-source models, our experimentation extends to an advanced closed-source model, GPT-3.5 (gpt-3.5-turbo-1106) (OpenAI, 2023), which is renowned for its superior performance in general capabilities, including safety.
@@
We compare our $\mathbb{IA}$ with five recently proposed defense methods, i.e., BPE-dropout (Jain et al., 2023), ICD (Wei et al., 2023), Self Defense (Helbling et al., 2023), Moral Self-Correction (Ganguli et al., 2022), and System-Mode Self-Reminder (Xie et al., 2023). Among them, the first two methods pre-process user queries to undermine harm while the last three exploit LLMs’ intrinsic capabilities of self-correction and improvement against jailbreak attacks. For a fair comparison, we closely follow the best default parameters in their papers.
@@
For the DAN dataset, we compile an evaluation dataset of 1560 samples by extracting 195 questions from each jailbreak community within the forbidden question set (Shen et al., 2023). For AdvBench, we follow Zou et al. (2023) and conduct transfer attacks on Vicuna-7B and 13B. The adversarial suffix achieving the lowest loss after 500 steps of optimization are adopted to further attack target models on 100 individual harmful behaviors (Wei et al., 2023). For open-source models, we download them from HuggingFace [^5] https://huggingface.co/models. For closed-source models, we obtain the responses of GPT-3.5 via API calls. Throughout our experiments, we set a temperature of zero for deterministic outcomes Peng et al. (2023) and a generation length of 1024 tokens, employing default system prompt templates for each LLM if available.

### v3

@@ Abstract
Aligning large language models (LLMs) with human values, particularly in the face of complex and stealthy jailbreak attacks, presents a formidable challenge. In this study, we present a simple yet highly effective defense strategy, i.e., Intention Analysis ($\mathbb{IA}$). The principle behind this is to trigger LLMs’ inherent self-correct and improve ability through a two-stage process: 1) essential intention analysis, and 2) policy-aligned response. Notably, $\mathbb{IA}$ is an inference-only method, thus could enhance the safety of LLMs without compromising their helpfulness [^1] Improving the safety of LLM with training-required methods Ouyang et al. (2022); Touvron et al. (2023) always necessitates great effort to strike a delicate balance between safety and helpfulness.. Extensive experiments on varying jailbreak benchmarks across ChatGLM, LLaMA2, Vicuna, MPT, DeepSeek, and GPT-3.5 show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -53.1% attack success rate) and maintain the general helpfulness. Encouragingly, with the help of our $\mathbb{IA}$, Vicuna-7B even outperforms GPT-3.5 in terms of attack success rate. Further analyses present some insights into how our method works. To facilitate reproducibility, we release our code and scripts at: https://github.com/alphadl/SafeLLM_with_IntentionAnalysis.
@@
Yuqi Zhang1, Liang Ding2, Lefei Zhang1 Dacheng Tao2 1School of Computer Science, Wuhan University 2School of Computer Science, The University of Sydney {zhangyuqi, zhanglefei}@whu.edu.cn, {liangding.liam, dacheng.tao}@gmail.com
@@ 1 Introduction
Figure 1: Performance of our method on different LLMs. Our $\mathbb{IA}$ 1) reduces Attack Success Rate ($\downarrow$) against both crafted jailbreak prompts (DAN and DeepInception) and automatic attack (GCG), 2) achieves remarkable safety improvements for both SFT (Vicuna-7B & MPT-30B-Chat) and RLHF (GPT-3.5) LLMs.
Despite numerous methods focusing on defending LLMs against these jailbreak attacks (Li et al., 2024; Jain et al., 2023; Chen et al., 2024), their efficacy markedly decreases when handling those with stealthy and complex intention (see experimental results in Section 4.2). This type of jailbreak attacks are typically hand-crafted prompts by using more instructions and employing specific techniques, such as creating a virtual scenario where LLMs are assigned the role of “DAN” (Do Anything Now, Liu et al., 2023b). Such complex and stealthy intents pose a great challenge for LLMs to maintain safety when confronting these jailbreak attacks (Shen et al., 2023; Bao et al., 2023).
@@
    We introduce $\mathbb{IA}$, a new method that significantly enhances LLM safety in the context of sophisticated jailbreak attacks through an intention analysis mechanism.
@@
    Empirically, our robust $\mathbb{IA}$ significantly and consistently reduces the harmfulness of LLM outputs, while maintaining the helpfulness, achieving new state-of-the-art performance on several benchmarks, e.g., DeepInception.
@@ 2 Related Work
Despite significant efforts to align LLMs with human preference (Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Korbak et al., 2023; Miao et al., 2024), adversarial attackers can still elicit harmful responses from LLMs by “jailbreak” attacks (Shen et al., 2023; Liu et al., 2023b). Current jailbreak attack methods are primarily classified into two categories: in-the-wild jailbreak prompts and optimization-based automatic attacks (Chao et al., 2023). In-the-wild jailbreak prompts are typically hand-crafted through human ingenuity and is semantically understandable in general Shen et al. (2023). For optimization-based automatic attacks, a representative work is to automatically fetch a transferable attack suffix through the Greedy Coordinate Gradient (GCG) algorithm which maximizes the probability of the language model generating an affirmative and unsafe response Zou et al. (2023).
Safety-Enhancing Defensive Methods
Recently, numerous methods have been developed to reduce LLMs’ harmful generations in inference stage. A branch of them mainly concentrates on controlling the content that LLMs can see by pre-processing the user inputs, such as perplexity filtering (Alon and Kamfonas, 2023; Jain et al., 2023), paraphrasing (Jain et al., 2023) and re-tokenization (Cao et al., 2023; Jain et al., 2023). Another branch focuses on exploiting LLMs’ intrinsic capabilities of self-correction and improvement against jailbreak attacks, such as letting LLMs self-evaluate their outputs (Helbling et al., 2023) or reminding of safety in system mode with conventional decoding Xie et al. (2023) or contrastive decoding Zhong et al. (2024).
While these existing methods show effectiveness in preventing LLMs from generating unsafe responses, their efficacy significantly diminishes when confronted with sophisticated jailbreak attacks due to the ignorance of the underlying stealthy and malicious intention. In contrast, our method handles such jailbreak attacks by leveraging the intrinsic intent recognition capabilities of LLMs, which leads to a significant enhancement of LLM safety (see Table 1 for details).
@@ 3.2 $\mathbb{IA}$: Intention Analysis
To achieve the above goal, we introduce $\mathbb{IA}$, a zero-shot intention analysis mechanism, to guide LLMs to explicitly identify and understand the underlying intention of a user query before facilitate a final response. Specifically, we devise a two-stage intention analysis instruction to accomplish the whole process [^2] Full prompts can be found in Figure 8., as illustrated in Figure 2(b): (1) essential intention analysis and (2) policy-aligned response.
@@
Models Defense Methods Jailbreak Methods Average
DAN SAP200 DeepInception GCG AutoDAN
ChatGLM-6B Vanilla 29.0 45.8 100 88.0 99.5 72.5
ICD 19.1 2.81 17.1 17.0 2.00 11.6
Self-Reminder 22.5 3.13 17.9 0.00 66.0 21.9
SmoothLLM 7.19 20.6 84.5 1.00 84.0 39.5
$\mathbb{IA}$ (Ours) 5.48 6.12 0.00 1.00 2.00 2.92
@@
ICD 0.98 0.00 0.00 0.00 0.00 0.20
Self-Reminder 0.77 0.00 4.38 0.00 0.00 1.03
SmoothLLM 0.31 2.81 86.5 0.00 71.5 32.2
$\mathbb{IA}$ (Ours) 0.13 0.00 0.00 0.00 0.00 0.03
@@
ICD 40.4 32.8 0.00 1.00 88.0 32.4
Self-Reminder 41.3 33.8 55.4 11.0 98.5 48.0
SmoothLLM 13.5 54.4 96.4 8.00 98.5 54.2
$\mathbb{IA}$ (Ours) 3.42 0.31 0.00 0.00 10.5 2.85
@@
ICD 53.9 32.8 86.9 0.00 91.5 53.0
Self-Reminder 52.5 36.9 75.7 1.00 83.0 49.8
SmoothLLM 17.3 37.0 94.0 5.00 98.0 50.3
$\mathbb{IA}$ (Ours) 0.94 1.12 0.00 0.00 3.50 1.11
@@
ICD 49.4 29.9 100 3.00 45.6
Self-Reminder 46.9 39.4 100 19.0 51.3
SmoothLLM 60.6 64.4 22.0 22.0 42.3
$\mathbb{IA}$ (Ours) 5.38 19.2 4.78 4.00 8.34
@@
ICD 45.6 14.4 47.8 9.00 29.2
Self-Reminder 9.58 7.81 3.19 1.00 5.40
SmoothLLM 26.9 11.9 51.0 0.00 22.4
$\mathbb{IA}$ (Ours) 3.78 1.56 7.57 2.00 3.73
@@
ICD 0.94 0.31 0.00 0.00 0.31
Self-Reminder 2.81 0.31 0.00 0.00 0.78
SmoothLLM 0.64 0.00 0.00 0.00 0.16
$\mathbb{IA}$ (Ours) 0.64 0.00 0.00 0.00 0.16
Table 1: Comparison of our $\mathbb{IA}$ and four baselines under five jailbreak methods in terms of ASR (%). The best average results are highlighted in bold. Among them, DAN, SAP200, and DeepInception are complex and stealthy in-the-wild jailbreaks, while GCG and AutoDAN are optimization-based automatic jailbreaks. “—” means lacking official AutoDAN implementation for distributed larger models (MPT-30B and DeekSeek-67B) or white-box LLM weights required (GPT-3.5).
@@ 4.1 Setup
For safety datasets, we experiment on three main categories of jailbreak attacks, including three representative complex and stealthy in-the-wild jailbreak datasets (i.e. DAN (Shen et al., 2023), SAP200 (Deng et al., 2023a), and DeepInception (Li et al., 2023)), two popular optimization-based automatic jailbreak methods (i.e. GCG (Zou et al., 2023) and AutoDAN (Liu et al., 2023a)), and two advanced attacks for GPT-3.5 (i.e. multilingual attack called MultiJail (Deng et al., 2023b) and encryption-based attack named CipherChat (Yuan et al., 2024)). Besides, to evaluate the effect of our $\mathbb{IA}$ on helpfulness for general harmless in-distribution queries, we conduct experiments on three widely recognized datasets, i.e., AlpacaEval (Dubois et al., 2023), MMLU (Hendrycks et al., 2021) and TruthfulQA (Lin et al., 2022).
Figure 3: The MultiJail (under two scenarios) and CipherChat Datasets results on GPT-3.5 with and without our $\mathbb{IA}$. (a) Results on direct MultiJail dataset including Chinese (zh), Italic (it), Vietnamese (vi), Arabic (ar), Korean (ko), Thai (th), Bengali (bn), Swahili (sw), and Javanese (jv). (b) Results on malicious jailbreak prompt attached to MultiJail. (c) Results on CipherChat including ASCII (en), UTF (zh), Unicode (zh), and SelfCipher (en and zh) encryptions.
@@
For safety assessment, we annotate the harmfulness of responses and report attack success rate (ASR, Shen et al., 2023), where lower scores indicate stronger safety. Specifically, for DAN dataset, considering the complexity of responses, we adopt gpt-3.5-turbo-0613555https://openai.com/blog/chatgpt as the auto-annotation function following Deng et al. (2023a) and carry our human evaluation in Appendix C.1 to ensure the credibility. For the rest safety datasets, we annotate harmfulness following Zou et al. (2023) by matching refusal strings (e.g., “I’m sorry”; see Appendix C.2 for detailed settings).
For helpfulness assessment, we report win rate (Dubois et al., 2023) for AlpacaEval and accuracy (Hendrycks et al., 2021) for MMLU. For TruthfulQA, we follow Chuang et al. (2023) and report on two distinct metrics: MC1 and MC2 scores, where higher scores indicate stronger factuality (see Appendix C.3 for more details).
@@
To evaluate the effectiveness of our $\mathbb{IA}$ method, we experiment on models varying scales and levels of alignment, including not only SFT models, i.e. Vicuna-7B/13B-v1.1 (Chiang et al., 2023) and MPT-30B-Chat (Team, 2023), but also RLHF models, i.e. ChatGLM-6B (Zeng et al., 2023), LLaMA2-7B-Chat (Touvron et al., 2023), and DeepSeek-67B-Chat (DeepSeek-AI, 2024). Beyond these representative open-source models, our experimentation extends to an advanced closed-source model, GPT-3.5 (gpt-3.5-turbo-1106) (OpenAI, 2023), which is renowned for its superior capabilities, including safety.
@@
We compare our $\mathbb{IA}$ with vanilla LLMs (without defense) and six popular defense methods, i.e., ICD (Wei et al., 2023b), (System-Mode) Self-Reminder (Xie et al., 2023), SmoothLLM (Robey et al., 2023), BPE-dropout (Jain et al., 2023), Self Defense (Helbling et al., 2023), and Moral Self-Correction (Ganguli et al., 2022). Among them, we report the first three representative defense methods in Table 1 and the entire in Table 9. For a fair comparison, we closely follow the best default parameters in their papers.
@@
The detailed $\mathbb{IA}$ prompts for experiments are provided in Figure 8666To assess the resilience of our method against specific expressions, we construct alternative $\mathbb{IA}$ prompts and conduct comparison experiment in Section 5.2.. For the DAN dataset, we compile an evaluation dataset of 1560 samples by extracting 195 questions from each jailbreak community within the forbidden question set (Shen et al., 2023). For GCG, we follow Zou et al. (2023) and conduct transfer attacks on Vicuna-7B and 13B. The adversarial suffix achieving the lowest loss after 500 steps of optimization are adopted to further attack target models on 100 individual harmful behaviors (Wei et al., 2023b). For open-source models, we download them from HuggingFace [^7] https://huggingface.co/models. For closed-source models, we obtain the responses of GPT-3.5 via API calls. Throughout our experiments, we set a temperature of zero for deterministic outcomes Peng et al. (2023) and a generation length of 1024 tokens, employing default system prompt templates for each LLM regarding their official reports.